# Entanglement, Trace Anomaly, and Confinement in QCD

### v3

@@ Acknowledgements.
Appendix
@@
Derivation of Eq. (1).— We work on the $t=0$ hypersurface and write $x^{\mu}=(t,r_{\parallel},\bm{r}_{\perp})$ with $r_{\parallel}\equiv x^{1}\in\mathbb{R}$ the longitudinal coordinate and $\bm{r}_{\perp}\equiv(x^{2},x^{3})\in\mathbb{R}^{2}$ the transverse vector ($\bm{x}\equiv(r_{\parallel},\bm{r}_{\perp})$). Its magnitude is $r_{\perp}=|\bm{r}_{\perp}|=b_{\perp}$ which can be identified with the impact-parameter space, and the three–dimensional radius is $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$. The volume element factorizes into longitudinal and transverse parts,
@@
and, with the substitutions $r_{\perp}=r\sin\theta$ and $r_{\parallel}=r\cos\theta$, this becomes the familiar spherical‑coordinate form
@@
1. Modular Hamiltonian after tracing out the exterior. Tracing over all quark‑ and gluon‑field modes with support outside a sphere of radius $R$, as in Fig. S1, produces the reduced state $\hat{\rho}_{<}(R)=\mathrm{Tr}_{r>R}|0\rangle\!\langle 0|$. Its generator of modular flow, $K(R)\equiv-\ln\hat{\rho}_{<}(R)$, governs time evolution for an observer confined to $r<R$. For the planar entangling surface that bounds the right Rindler wedge $\mathcal{R}^{+}=\{x^{1}>|t|\}$[6], the Bisognano–Wichmann theorem [91, 92] identifies the modular Hamiltonian with the boost generator,
@@
Because $K$ couples linearly to $T_{00}$, the eigenvalue density of $\hat{\rho}_{<}$ (the entanglement spectrum) is fixed by $\langle T_{00}\rangle$. In the Lorentz‑invariant vacuum with the mostly‑plus Minkowski metric, $T_{00}=-\tfrac{1}{4}T^{\mu}{}_{\mu}$; tracing Lorentz indices therefore ties the spectrum directly to the trace spectral density $\rho_{h}$.
2. Conformal map to a ball. A special conformal transformation with $a^{\mu}=(0,0,0,1/R)$ maps the half‑space $r_{\parallel}>0$ to the ball $r<R$ and rescales the spherical measure $r^{2}\sin\theta$:
@@
Because $T_{00}$ has conformal weight 4, this Jacobian cancels and one obtains the sphere‑local modular Hamiltonian [13, 14]
@@
4. Surface–local two‑point function. Inserting (S6) into (S5) gives
@@
The bulk contribution integrates to a power divergence (e.g. area‑law $\propto R^{2}/\epsilon^{2}$, curvature terms on $\Sigma_{\perp}$, etc.) removable by counter‑terms and therefore does not affect the universal logarithmic coefficient. The coefficient of $\delta(R-r)$ is unity by construction (Gauss’ theorem). Using (S8) in (S7) and keeping only the surface term,
@@
5. Entanglement‑entropy flow. A uniform Weyl shift $\delta g_{\mu\nu}=2\sigma g_{\mu\nu}$ gives [2]
@@
where we used (S9) to collapse the 3D integral to 2D. Choosing $\sigma=\delta R/R$ gives the RG‐flow equation
@@
Integrating from $\epsilon$ to $R$, defining $J(R^{\prime})=\int_{r=R^{\prime}}\mathrm{d}^{2}r_{\perp}\,T(0;r_{\parallel},\bm{r}_{\perp})$, and discarding the analytic (counter‑term) piece yields the universal logarithm:
@@
The slope of the Lorentz–invariant scalar form factor $A^{S}_{q,g}(t)\equiv\eta_{\mu\nu}T^{\mu\nu}_{q,g}(t)$ at the origin defines a frame–independent rms radius, while the slope of the energy density $A^{00}_{q,g}(t)=T^{00}_{q,g}(t)$—taken in the nucleon rest frame—defines a distinct Breit–frame radius:

### v4

@@ Acknowledgements.
DATA AVAILABILITY
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Supplemental Material
@@
Derivation of Eq. (1).— We work on the $t=0$ hypersurface and write $x^{\mu}=(t,r_{\parallel},\bm{r}_{\perp})$ with $r_{\parallel}\equiv x^{1}\in\mathbb{R}$ the longitudinal coordinate and $\bm{r}_{\perp}\equiv(x^{2},x^{3})\in\mathbb{R}^{2}$ the transverse vector ($\bm{x}\equiv(r_{\parallel},\bm{r}_{\perp})$). Its magnitude is $r_{\perp}=|\bm{r}_{\perp}|=b_{\perp}$ which can be identified with the impact-parameter space, and the three-dimensional radius is $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$. The volume element factorizes into longitudinal and transverse parts,
@@
and, with the substitutions $r_{\perp}=r\sin\theta$ and $r_{\parallel}=r\cos\theta$, this becomes the familiar spherical-coordinate form
@@
1. Modular Hamiltonian after tracing out the exterior. Tracing over all quark- and gluon-field modes with support outside a sphere of radius $R$, as in Fig. S1, produces the reduced state $\hat{\rho}_{<}(R)=\mathrm{Tr}_{r>R}|0\rangle\!\langle 0|$. Its generator of modular flow, $K(R)\equiv-\ln\hat{\rho}_{<}(R)$, governs time evolution for an observer confined to $r<R$. For the planar entangling surface that bounds the right Rindler wedge $\mathcal{R}^{+}=\{x^{1}>|t|\}$[6], the Bisognano–Wichmann theorem [91, 92] identifies the modular Hamiltonian with the boost generator,
@@
Because $K$ couples linearly to $T_{00}$, the eigenvalue density of $\hat{\rho}_{<}$ (the entanglement spectrum) is fixed by $\langle T_{00}\rangle$. In the Lorentz-invariant vacuum with the mostly-plus Minkowski metric, $T_{00}=-\tfrac{1}{4}T^{\mu}{}_{\mu}$; tracing Lorentz indices therefore ties the spectrum directly to the trace spectral density $\rho_{h}$.
2. Conformal map to a ball. A special conformal transformation with $a^{\mu}=(0,0,0,1/R)$ maps the half-space $r_{\parallel}>0$ to the ball $r<R$ and rescales the spherical measure $r^{2}\sin\theta$:
@@
Because $T_{00}$ has conformal weight 4, this Jacobian cancels and one obtains the sphere-local modular Hamiltonian [13, 14]
@@
4. Surface-local two-point function. Inserting (S6) into (S5) gives
@@
The bulk contribution integrates to a power divergence (e.g. area-law $\propto R^{2}/\epsilon^{2}$, curvature terms on $\Sigma_{\perp}$, etc.) removable by counter-terms and therefore does not affect the universal logarithmic coefficient. The coefficient of $\delta(R-r)$ is unity by construction (Gauss’ theorem). Using (S8) in (S7) and keeping only the surface term,
@@
5. Entanglement-entropy flow. A uniform Weyl shift $\delta g_{\mu\nu}=2\sigma g_{\mu\nu}$ gives [2]
@@
where we used (S9) to collapse the 3D integral to 2D. Choosing $\sigma=\delta R/R$ gives the RG-flow equation
@@
Integrating from $\epsilon$ to $R$, and discarding the analytic (counter-term) piece yields the universal logarithm:
@@
The slope of the Lorentz-invariant scalar form factor $A^{S}_{q,g}(t)\equiv\eta_{\mu\nu}T^{\mu\nu}_{q,g}(t)$ at the origin defines a frame-independent rms radius, while the slope of the energy density $A^{00}_{q,g}(t)=T^{00}_{q,g}(t)$—taken in the nucleon rest frame—defines a distinct Breit-frame radius: